OpenStudy (anonymous):

Suppose that the trace of a 2×2 matrix A is tr(A)=3, and the determinant is det(A)=−40. Find the eigenvalues of A

ganeshie8 (ganeshie8):

I think you need to solve below : \(\large \lambda^2 - \lambda (\text{trace} ) + \text{determinant} = 0 \)

OpenStudy (anonymous):

\[ \begin{bmatrix} a&b\\ c&d \end{bmatrix} \]

OpenStudy (anonymous):

\[ (a-\lambda)(d-\lambda)-bc=0 \]We know \(ad-bc = -40\) and \(a+d=3\).

OpenStudy (anonymous):

Make that \(\lambda_1\) and \(\lambda_2\).

OpenStudy (anonymous):

\[ ad-(a+d)\lambda+\lambda^2-bc = 0 \]Then we get: \[ \lambda^2-(\color{red}{a+d})\lambda +(\color{blue}{ad-bc})=0 \]So looking at that, I agree with @ganeshie8

OpenStudy (anonymous):

Thank you! I got my small eigenvalue and large eigenvalues mixed up. small eigen.= -5, and large eigen. = 8